Given that sinA=1/2 , find the vlue of sinAcos(90degree-A).
im lost..

Answer by chibisan(131)

Given that sinA=1/2 , find the vlue of sinAcos(90degree-A)

cos(90degree-A) = sinA
so
sinA*sinA = 1/2*1/2 = 1/4